OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y+1=4(x+3) y+1=4x+12 y=4x+11